OpenStudy (anonymous):

Lines a and b are parallel and m<2=102 degrees What is m<5?

OpenStudy (anonymous):

m<6 = m<2 = 102 degrees (corresponding angles). m<6 and m<5 are complementary angles (means they add to 180 degrees). m<6 + m<5 = 180 m<5 = 180 - m<6 = 180 - 102 = ?

OpenStudy (anonymous):

78?

OpenStudy (anonymous):

@barreraA

OpenStudy (barreraa):

so what do u think

OpenStudy (anonymous):

idk i think its 78 degrees